Need help:
In how many ways can 4 letters be chosen from the set {a, b, c, d, e, f} if order is important and no repeats are allowed?
A) 15
B) 1296
C) 360
D) 24

Answer by user_dude2008(1862)

6 P 4 = 6!/(6-4)! = 6!/2! = 720/2 = 360

Ans: C) 360